OpenStudy (seattle12345):

Can some one help with this?: If f(x)= 10^x, show that f(x+h)-f(x)/h = 10^x(10h-1/h)

OpenStudy (anonymous):

\[f(x)= 10^x \] \[f(x+h)= 10^{x+h}=10^x10^h\] just substitute those into the given equation\[\frac{f(x+h)-f(x)}{h }\] and simplify

OpenStudy (seattle12345):

so would it be \[\frac{ 10^{x}+10^{h}-10^{x}}{ h }\] and the factor out \[10^{x}\]?

OpenStudy (seattle12345):

@completeidiot

OpenStudy (anonymous):

f(x+h) was multiplying not adding them

OpenStudy (anonymous):

\[10^x10^h \neq 10^x+10^h\]

OpenStudy (anonymous):

it would come out to be \[\frac{10^x10^h - 10^x}{h}\] and like you said, factor out the \[10^x\]

OpenStudy (seattle12345):

@completeidiot oh ok, so this is probably a dumb question but is \[10^{xh}=10^{x}10^{h}\]

OpenStudy (anonymous):

not a dumb question no it is not \[10^{xh}=(10^x)^h=(10^h)^x\] \[10^{x+h}=10^x10^h\] i think i wrote that right im quite sure i did

OpenStudy (anonymous):

http://www.purplemath.com/modules/exponent.htm

OpenStudy (seattle12345):

thanks im going to need more practice lol

OpenStudy (anonymous):

no prob